1) Loan amortization and EAR You want to buy a car, and a local bank will lend you $10,000. The loan will be fully amortized over 5 years (60 months), and the nominal interest rate will be 11% with interest paid monthly. What will be the monthly loan payment? Round your answer to the nearest cent. $ What will be the loan's EAR? Round your answer to two decimal places. %

2)

Effective rate of interest

Find the interest rates earned on each of the following. Round each answer to two decimal places.

  1. You borrow $710 and promise to pay back $781 at the end of 1 year.

    %

  2. You lend $710 and the borrower promises to pay you $781 at the end of 1 year.

    %

  3. You borrow $86,000 and promise to pay back $370,698 at the end of 14 years.

    %

  4. You borrow $20,000 and promise to make payments of $6,687.60 at the end of each year for 5 years.

    %
Answer #1

1

PVOrdinary Annuity = C*[(1-(1+i/(f*100))^(-n*f))/(i/(f*100))]
C = Cash flow per period
i = interest rate
n = number of payments I f = frequency of payment
10000= Cash Flow*((1-(1+ 11/1200)^(-5*12))/(11/1200))
Cash Flow = 217.42
Using Calculator: press buttons "2ND"+"FV" then assign
PV =-10000
I/Y =11/12
N =5*12
FV = 0
CPT PMT
Using Excel
=PMT(rate,nper,pv,fv,type)
=PMT(11/(12*100),12*5,,10000,)
EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
? = ((1+11/(12*100))^12-1)*100
Effective Annual Rate% = 11.57
Please ask remaining parts seperately, questions are unrelated